OpenStudy (anonymous):

A grid shows the positions of a subway stop and your house.The subway stop is located at (-6,1) and your house is located at (-2.-8).What is the distance to the nearest unit between your house and the subway stop? 10 20 15 7

OpenStudy (anonymous):

If you just calculate the shortest distance: c^2 = sqrroot(a^2 + b^2) where a = -5-0 b = 2 - (-7) so c^2 = sqrroot((-5)^2 + 9^2) c^2 = 25 + 81 = 106 c = 10